3. We discussed in class the fact that when it comes to calculating the Annual Worth (AW) of alternatives have different lives. Why the AW calculations and analyses are done this way compared to those of the Present Worth (PW) and the Future worth (FW) analyses alternatives, we bse the actual life (n) for each alternative even if the and calculations? 4. A small company would like to set aside the amount of $90,000 now to purchase a new piece of equipment after 5 years from now that cost of $200.000. will the company be able to purchase the equipment at an annual interest rate of 6%? 5. Compare the two economic altermatives below and select the best one based on their in st nie of 6% for b shrm ines Allal Annual worth Use an annual i Aternative Initial Cost Yearly C Value ate tyears
Answer #1

3. Annual worth is comparison of annual value of two or more alternatives even if they have different span of life or economic life. But present worth and future worth of alternatives provide specific value of alternatives that have same or varying span of life. Both present worth and future worth analysis are used in comparison of projects, plant machinery and equipment. But annual worth analysis is mostly suited for assessing the worthiness of plant machinery and marketing or business plans of a company.

4) rate of interest = 6% period =5 years

Initial sum of the company for machine purchase = $90000

therefore, the amount of money in 5 years = 90000(1+.06)5 = 120440.30

Also cost of the equipment after 5 years = $200000

therefore, the company will not be able to purchase the equipment after 5 years

5) Annual worth of alternative M (AWM ) = -30000(A/P, 6%,5)-6000(A/F, 6%,5) +8000(A/P, 6%,5) +900(A/F,6%,5)

                                                    = -30000*.2374-6000* .1774 +8000*.2374+900*.1774 =$-6127.54

Annual worth of alternative N (AWN ) = -45000(A/P, 6%,2) -4500(A/F, 6%,2)+6500(A/P,6%,2)

                                          = -45000* .5454 -4500*.4854+6500*.5454 =$-23182.2

As AWM > AWN therefore, alternative M is best alternative.